OpenStudy (help2001):

)A pair of equations is shown below. x + y = 2 y = one halfx + 5 If the two equations are graphed, at what point do the lines representing the two equations intersect? (4, −2) (−2, 4) (2, 5) (5, −2)

OpenStudy (breanna1999):

its the second choice

OpenStudy (help2001):

WHICH ONE!!!!!?

OpenStudy (anonymous):

its the second choice because none of the others make sense to the question

OpenStudy (help2001):

lol add some explaination guys i cant belive someone who just says the option and doesnt give an explaination

OpenStudy (breanna1999):

the second one. Plug in the x and y values to double check each equations

OpenStudy (help2001):

thats not an explaination @rock913335 an explaination is what @breanna1999 said

OpenStudy (help2001):

why is their only moms looking at my post lol

OpenStudy (anonymous):

sorry

OpenStudy (breanna1999):

im not a mom lmao. Im only 16

OpenStudy (anonymous):

x + y = 2 y = one halfx + 5 so 1. = (4, −2) 4+-2=2 but 5+2 doesnt equal -2 2. -2+4=2 which is good then -1+5 does equal 4 so if you can only choose one answer its number 2

OpenStudy (anonymous):

lololol same im 16